OpenStudy (anonymous):

What is the perimeter of a rectangle with length six and one over two inches and width 3 inches? 9 inches nine and one over two inches 16 inches 19 inches

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (math&ing001):

Perimeter = ( 6.5 + 3 ) * 2 = 9.5 * 2 = 19 inches